OpenStudy (ammarah):

Tell whether the parabola opens up, down, left ,or right x^2=6y -7x^2=y x^2=6y -7x^2=y

OpenStudy (ammarah):

i know that ifa prabola is pos it goes up and neg down but idk how to do left and right....

OpenStudy (ammarah):

@mathstudent55 @goformit100

OpenStudy (mathmale):

You can recognize a vertical parabola where you see y to the first power and x to the 2nd. Similarly, you can recognize a horiz. parabola where you see x to the first power and y to the second. So, does x^2=6y open to the left or to the right up or up, or down?

OpenStudy (ammarah):

ok so right?

OpenStudy (mathmale):

Mind answering my question? Thanks.

OpenStudy (ammarah):

yeah i said right

OpenStudy (ammarah):

? heloo

OpenStudy (mathmale):

You meant that your parabola opens TO the right? "yeah i said right"

OpenStudy (mathmale):

You've typed out four different answer choices. Does that mean, to you at least, that only one of them is correct? Was there an illustration with this problem, or just the problem stated in words?

OpenStudy (ammarah):

words

OpenStudy (ammarah):

wait i dont understand you so the fist one is correct?

OpenStudy (ammarah):

The second one would be left? The third is right again the last one is left idk why they did it again

OpenStudy (mathmale):

Something's not quite right here. You supposedly have four possible answers, but your first and third appear to be the same, and your second and four also appear to be the same. That leaves us with two pairs of identical equations. ??? x^2=6y -7x^2=y x^2=6y -7x^2=y

OpenStudy (ammarah):

im sorry i think on the hw there was a typo i only want to know if two of the answers are correct so x^=6y is right and -7x^=y is left am i correct?

OpenStudy (mathmale):

Looking at the first equation, I see x is squared and y is not squared. Which of the two cases have you here? I mentioned the correct criteria earlier: You can recognize a vertical parabola where you see y to the first power and x to the 2nd. Similarly, you can recognize a horiz. parabola where you see x to the first power and y to the second. " x^=6y is right "? No. I think you meant " x^2 = 6 y". But no, that does not open to the right. See the criteria, above. " -7x^=y " ? again, but sorry, no. Something is missing from this equation; can you find what's missing and write it in?

OpenStudy (ammarah):

the x ^2 and so the fist one is left and other one is right

OpenStudy (ammarah):

the secone one is down

OpenStudy (mathmale):

My response: 'You can recognize a vertical parabola where you see y to the first power and x to the 2nd." You have x^2 in all four of the possible answer choices (unless you have a typo or two there); thus, ALL of these parabolas open either UP or DOWN and none open left or right.

OpenStudy (ammarah):

up, down im soo sorry i toook long on this it confused me but thanks i need more helpp

OpenStudy (mathmale):

Can't promise I'll be on OpenStudy much longer. I'd suggest y ou post your next problem in the usual way; I'll respond if I believe I can help you with it in the time I have left online.